A bag contains 9 red marbles, 8 white marbles, and 6 blue marbles. Randomly choose 2 marbles, one at a time, and without replacement. What is the probability that both are white?

Select one:

a. (8/23)*(7/23) = 56/529 = 0.106

b. (8/23)*(7/22) = 56/506 = 0.111

c. (8/23) + (7/22) = 337/506 = 0.6660

d. (1/8)*(1/7) = 1/56 = 0.018
Answer #1

Answer: b) (8/23)*(7/22) = 56/506 = 0.111

Explanation:

If we choose 2 marbles, one at a time and without replacement then the probability of first white marbles is (8/23) and the probability of second white marble is (7/22). Thus the probability that both are white is =(8/23)*(7/22) = 56/506 = 0.111